OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form. Answer 4x + y = 7 4x - y = 3 4x - y = -22

OpenStudy (anonymous):

y - 2 = 4(x + 5) y - 2 = 4x + 20 y = 4x + 20 + 2 y = 4x + 22

OpenStudy (anonymous):